Telescope.org
This website uses cookies. To read more about our use of cookies, click here.
To not be bothered by this message again, click the OK button.
OK

Current Status

The problems with original BRT hardware were not repairable. This telescope will now be decommissioned.

In its place, our new observatories are now linked with telescope.org are currently being tested. An announcement will be made when this is ready for general use. If you would like to get involved with testing, please drop us an email.

Current known issues:
  • The filter wheel sometimes reports the wrong filter
  • The camera attachment has come loose and this is causing occasional, random field rotation between images
  • The camera bias values are sometimes changing and this is causing calibration issues. This is a camera hardware fault.

A servicing trip is being arrange to fix these problems.

Planned Outages

None


Last updated: Thursday 30 March 2017, 12:03



Back to the index page

© 2016 The Open University